# Disorder Suppression of Charge Density Wave in the Honeycomb Holstein Model

### v1

@@ Abstract
This study investigates the effect of disorder on the CDW order in the Holstein model on the honeycomb lattice using DQMC simulations. Disorder is introduced via random hopping amplitudes $t_{ij}$, which are uniformly distributed around a mean value. The primary finding is that disorder suppresses the CDW phase, as evidenced by a reduction in the CDW structure factor $S_{\text{CDW}}$ and an increase in the critical inverse temperature $\beta_{c}$, indicating that stronger disorder stabilizes the CDW phase against thermal fluctuations. The analysis of charge-density correlations and finite-size scaling reveals that moderate to strong disorder disrupts long-range CDW order through Anderson localization, leading to the failure of conventional critical scaling. Additionally, disorder suppresses electron kinetic energy and dc conductivity, further impairing electronic transport. Spectral function calculations show that disorder broadens and eventually closes the CDW energy gap. These results provide important insights into the stability and critical behavior of correlated electronic phases in disordered 2D Dirac systems.
V Acknowledgement
@@
G.C.L and L.F.Z were supported by Key Laboratory of Multiscale Spin Physics with Contract No. SPIN2024N04 and “the Fundamental Research Funds for the Central Universities” with Contract No. 2024ZCJH13. We acknowledge computational support from the HSCC of Beijing University of Posts and Telecommunications. This work is supported by NSFC (12474218) and Beijing Natural Science Foundation (No. 1242022 and 1252022).

### v2

@@ IV Summary
This study investigates the effect of disorder on the CDW order in the Holstein model on the honeycomb lattice using DQMC simulations. Disorder is introduced via random hopping amplitudes $t_{ij}$, which are uniformly distributed around a mean value. The primary finding is that disorder suppresses the CDW phase, as evidenced by a reduction in the CDW structure factor $S_{\text{CDW}}$ and an increase in the critical inverse temperature $\beta_{c}$, indicating that stronger disorder destabilizes the CDW phase and renders it more susceptible to thermal fluctuations. The analysis of charge-density correlations and finite-size scaling reveals that moderate to strong disorder disrupts long-range CDW order through Anderson localization, leading to the failure of conventional critical scaling. Additionally, disorder suppresses electron kinetic energy and dc conductivity, further impairing electronic transport. Spectral function calculations show that disorder broadens and eventually closes the CDW energy gap. These results provide important insights into the stability and critical behavior of correlated electronic phases in disordered 2D Dirac systems.
Acknowledgements.
@@
G.C.L and L.F.Z were supported by the Fundamental Research Funds for the Central Universities from the Beijing University of Posts and Telecommunications under Grant No. 2024ZCJH13 and 2025JCTP08. This work was supported by NSFC (12474218) and Beijing Natural Science Foundation (No. 1242022 and 1252022). We also supported by The Key Laboratory of Multiscale Spin Physics with Contract No. SPIN2024N04. We acknowledge computational support from the HSCC of Beijing University of Posts and Telecommunications.
Appendix A FINITE-SIZE SCALING ANALYSES FOR DIFFERENT ELECTRON-PHONON COUPLING STRENGTHS
Figure 8: The crossing plots and best data collapse of $S_{\rm CDW}/L^{7/4}$ for different disorder strengths $\Delta$. (a), (c), (e), (g) show the variation of $S_{\rm CDW}/L^{7/4}$ with $\beta$ for lattice sizes $L=4,5,6$; (b), (d), (f), (h) present the corresponding data collapse (via 2D Ising critical exponents) as a function of $L(\beta-\beta_{c})$. Here, $\lambda_{\rm D}=3/8$ and $\omega_{0}/t=1$.
Figure 9: The crossing plots and best data collapse of $S_{\rm CDW}/L^{7/4}$ for different disorder strengths $\Delta$. Here, $\lambda_{\rm D}=27/50$ and $\omega_{0}/t=1$.
Figure 10: The crossing plots and best data collapse of $S_{\rm CDW}/L^{7/4}$ for different disorder strengths $\Delta$. Here, $\lambda_{\rm D}=24/25$ and $\omega_{0}/t=1$.
Figure 11: The crossing plots and best data collapse of $S_{\rm CDW}/L^{7/4}$ for different disorder strengths $\Delta$. Here, $\lambda_{\rm D}=121/150$ and $\omega_{0}/t=1$.
This appendix (including Fig. 8, Fig. 9, Fig. 10 and Fig. 11) presents supplementary finite-size scaling results for the CDW structure factor $S_{\mathrm{CDW}}$ in systems with increased electron-phonon coupling. We examine four cases corresponding to the original coupling parameters $\lambda=1.5,\;1.8,\;2.2,\;2.4$, which translate to the following dimensionless coupling strengths: $\lambda_{D}=\lambda^{2}/(\omega_{0}^{2}W)=3/8,\;27/50,\;121/150,\;24/25$, respectively. All analyses employ the critical exponents $\gamma=7/4$ and $\nu=1$ of the two-dimensional Ising model [6], consistent with the main text. The fixed parameters throughout are $\omega_{0}/t=1$, half-filling $\rho=1$, lattice sizes $L=4,5,6$, and disorder strengths $\Delta=0.0,\;0.5,\;1.0,\;1.5$. These results further clarify the combined influence of EPC and disorder on the CDW critical behavior, extending the analysis presented in the main text for $\lambda_{D}=2/3$.
Appendix B S-WAVE PAIRING SUSCEPTIBILITY ANALYSIS
Figure 12: The s-wave pairing susceptibility $\chi_{s}$ as a function of inverse temperature $\beta$ for $L=4$, $\lambda_{D}=2/3$, and disorder strengths $\Delta=0.0,\;0.5,\;1.0,\;1.5$.
To investigate superconducting pairing correlations in the disordered honeycomb Holstein model, we compute the s-wave pairing susceptibility $\chi_{s}$, a key quantity for characterizing s-wave superconducting correlations in two-dimensional systems. It is defined as
$\chi_{s}=\frac{1}{N}\int_{0}^{\beta}d\tau\,\bigl\langle\Delta(\tau)\Delta^{\dagger}(0)\bigr\rangle,$
where $\Delta(\tau)=\sum_{i}\hat{d}_{i\uparrow}(\tau)\hat{d}_{i\downarrow}(\tau)$ is the imaginary-time dependent s-wave pairing operator, and $\hat{d}_{i\sigma}(\tau)=e^{\mathcal{H}\tau}\hat{d}_{i\sigma}e^{-\mathcal{H}\tau}$ denotes the imaginary-time evolved electron annihilation operator at site $i$ with spin $\sigma$.
All calculations use the same core parameters as the main text: lattice size $L=4$, dimensionless electron–phonon coupling $\lambda_{D}=2/3$, phonon frequency ratio $\omega_{0}/t=1$, and half-filling $\rho=1$. Disorder strengths $\Delta=0.0,\;0.5,\;1.0,\;1.5$ are varied to examine the influence of disorder on s-wave pairing correlations.
Fig. 12 displays the s-wave pairing susceptibility $\chi_{s}$ as a function of $\beta$ for different disorder strengths $\Delta$. The key result is that the $\chi_{s}$ curves for $\Delta=0.0$, $0.5$, $1.0$, and $1.5$ nearly coincide across the entire temperature range. This overlap indicates that the temperature dependence of $\chi_{s}$ is virtually identical regardless of the disorder strength. Most importantly, even at the lowest temperatures ($\beta\gtrsim 9$), there is no measurable difference between the susceptibility of the clean system and that of the strongly disordered system ($\Delta=1.5$).
This result clearly demonstrates that disorder has no significant impact on the s-wave pairing susceptibility in this system. This finding complements the main-text conclusion that disorder strongly suppresses CDW correlations, highlighting the selective role of disorder: it disrupts charge order while leaving s-wave pairing correlations essentially unaffected in the honeycomb Holstein model.